light of wavelength 450 nm is incident on a diffraction grating on which 5000 lines can ruled out determine the angle corresponding to third order


1
Expert's answer
2021-09-20T15:28:56-0400

"5000 \;lines/cm = 500000 \;lines/m \\

d = \frac{1}{500000}=2 \times 10^{-6} \\

d \times sin θ = m \times λ"

for third order m = 3

"sin θ = \frac{3λ }{d} \\

sin θ = \frac{3 \times 450 \times 10^{-9}}{2 \times 10^{-6}} \\

sin θ = 0.675 \\

θ = 71.67°"